# Large $N$ limit of Wilson loops on orientable closed surfaces in the light of Koike-Schur-Weyl duality and Spin Networks

### 11 Mar 2026 v1

@@ 4.5.2. Upper-bound II: bound for shortest words representatives
when $\tau_{u},\tau_{b}$ and $\varphi=\pi_{\alpha,\alpha}$ are respectively compatible permutation diagrams and an admissible quadrangle diagram, hence with $\tau_{u},\tau_{b}\in\mathcal{D}^{0}_{\mathfrak{a},\mathfrak{b}},\alpha\in S_{\mathfrak{a},\mathfrak{b}}.$ Under this assumption, let us rename $\varphi$-arcs $\alpha$-arcs. For such a Brauer map each $\mathcal{C}_{II}$-face of $\Sigma_{\tau_{u},\pi,\tau_{b},\varphi}$ can be identified with a rectangle with two parallel $\alpha$-arcs and two parallel $\iota$-arcs labeled by a letter of $\{a_{1},b_{1},\ldots,b_{g}\}$ and its inverse. Recall that each connected component of $\partial\Sigma_{\tau_{u},\pi,\tau_{b},\varphi}$ is the concatenation of $\tau,\pi$ and $\iota$ intervals. Since $\tau_{u},\tau_{b},\pi,\mathrm{id}_{s}$ are compatible for $\mathfrak{b},$ vertices along each boundary component have same $\mathfrak{b}$-colouring say $B\in\mathcal{S}$. In the latter case call the boundary component $B$-boundary. Since $\tau_{u},\tau_{b}\in\mathcal{D}_{\mathfrak{a},\mathfrak{b}}^{0}$, reading the $\mathfrak{a}$-value of $\iota$ intervals along an oriented $B$-boundary spells out a power $p$ of a cyclic permutation of the word labeled $B$ for each cycle of length $p$ of $\tau.$ We then assign the multiplicity of the boundary to be $p.$ Note that $p=1$ when $B\in\{W,W^{-1}\},$ since, as permutations, $\tau_{u},\tau_{b}$ act trivially on $\mathfrak{b}^{-1}(\{W,W^{-1}\}).$ Counted with their multiplicity, there are $n$ boundaries coloured $R,$ $m$ coloured $R^{-1},$ one $W$ and one $W^{-1}.$ Recalling the constraint (148), since $\alpha\in S_{\mathfrak{a},\mathfrak{b}},$ an $\alpha$-arc cannot have one endpoint in a $R$-boundary and the other in a $R^{-1}$-boundary.
@@
We will next compute and bound $\chi(\Sigma_{\tau_{u},\tau_{b},\omega,\alpha})$ cutting $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$ along some $\alpha$-arcs, extracting discs or annuli along $W$ and $W^{-1}$-boundaries of $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$, such that each component has in its boundary only $W$-intervals or only $W^{-1}$ intervals. [^28] Boundary components of pieces shall be concatenation of $\alpha$-arcs and of boundary intervals. Since there are no $RR^{-1}$-arcs, boundary component of pieces shall also include only $R$-intervals, or only $R^{-1}$-intervals. As in [Mag25], for these ”good” components, called below pieces, the shortest length condition on $\omega$ allows to upper bound the number of $W^{\pm}R^{\pm},W^{\pm}R^{\mp}$ arcs by the ones of $RR$ or $R^{-1}R^{-1}$ arcs. [^29] Allowing components with $WW^{-1}$-arcs would break this bound. The main points here is then to prove that a variant of [Mag25] yields the correct Euler characteristic bound for the remaining components despite having possibly more types of components. The argument can be split in two, with a first part bounding the Euler characteristics, denoted $\chi_{-}$ below, of a ”neighbourhood” of the good components.
@@
Assume $X$ is a $W^{\varepsilon}R^{\varepsilon^{\prime}}$-pre-piece or a $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arc for some $\varepsilon,\varepsilon^{\prime}\in\{-1,1\}.$ Thanks to Lemma 4.16, when $X$ is a pre-piece, among pre-pieces and $WR^{*}$-arcs, $\overline{X}$ only intersects closures of $W^{\varepsilon}R^{\varepsilon^{\prime}}$-pre-pieces and $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs along the two $W^{\varepsilon}R^{\varepsilon^{\prime}}$ in its boundary. Consider a connected set $P^{\prime}$ union of pre-pieces and $WR^{*}$-arcs with $P^{\prime}\supset X.$ By induction on number of pre-pieces $\overline{P}^{\prime}$ intersects, $P^{\prime}$ is either an open annulus or a disc $D$ with two distinct $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs in its boundary, and $P^{\prime}$ only intersects arcs and pre-pieces of type $W^{\varepsilon}R^{\varepsilon^{\prime}}.$
@@
The side length $e(P)$ is the number of $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs endpoints included in $\partial\Sigma_{W^{\varepsilon}}\cap\overline{P}.$ Going throughout $\partial\Sigma_{W^{\varepsilon}}\cap\overline{P}$ according to its orientation and reading letters for each $\alpha$-arc crossed, spells out a word or a cyclic equivalence of a word $\omega_{P}$ when $\chi(P)$ is respectively $1$ or $0.$
@@
The proof is a variation of [MP23, Lemma 5.18]. For completeness, let us detail the argument in our notations.
Recall the Birman-Series bound. When $\omega_{1},\omega_{2}$ are two words with letters in $\mathcal{A},$ let us write $\omega_{1}\prec_{c}\omega_{2}$ when $\omega_{1}$ is a subword of a cyclic permutation of $\omega_{2},$ and $\omega_{1}\sim_{c}\omega_{2}$ when $\omega_{1}$ is cyclic permutation of $\omega_{2}.$ Consider the word
@@
Since it is the main place where the shortest length condition plays a role, let us sketch a proof for completeness.
@@
Since there is no long chain and no long bloc, $d_{m}\leq 1$ for all $m$ and whenever $d_{i}=1=d_{j}$ for some $1\leq i<j\leq l$, there is $i<k<j$ with $d_{k}=-1.$ Consequently $\sum_{i=1}^{m}d_{i}\leq 1$ and (172) follows from the last display. ∎
@@
Since all connected components of $\Sigma^{*}$ and $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$ have non-empty boundary, they deform retract to the dual graph associated to any cell decomposition. Denote $\hat{\mathcal{G}}$ the graph dual to the $1$-skeleton of $X_{\tau_{u},\tau_{b},\omega,\alpha}$, so that vertices are labeled by $2$-cells of $X_{\tau_{u},\tau_{b},\omega,\alpha}$ and edges by $\alpha$-arcs. Let $\hat{\mathcal{G}}_{RR}$ be the graph obtained from $\hat{\mathcal{G}}$ by removing all edges given by $RR$ and $R^{-1}R^{-1}$-arcs. Then $\Sigma^{*}$ deforms retract to $\hat{\mathcal{G}}_{R}$ so that
@@
The only contribution to $\chi(\Sigma^{*})$ therefore comes form joint discs. Let us split the sum (178) as
@@
Since type of boundaries alternate between $R^{*}$ and $W^{*}$ boundaries when going through $\partial D$, changing whenever crossing a $WR^{*}$-arc, $d_{WR^{*}}(D)$ is even. If $d_{WR^{*}}(D)\geq 4,$ then
@@
Since any $WR^{*}$-arc bounding a joint disc is also bounding a piece, while any piece boundary goes through exactly two distinct oriented $WR^{*}$-arcs,353535with the convention that for $WR^{*}$-arc that do not belong to any pre-piece, the boundary cycle is exactly the arc with some orientation followed by its reverse.
@@
Denote by $\mathfrak{he}^{\prime}_{RR^{*}}$ the number of oriented $RR^{*}$-arcs for which no piece boundary goes through positively. Since each oriented $RR^{*}$-arc is positively bounding at most one piece,
@@
If $D$ is a $R^{\varepsilon}$-disc, $d^{*}(D)=0$ and for each oriented interval of $\partial\Sigma_{R^{\varepsilon}}$ meeting its boundary, the $\mathfrak{a}$-values $x,y$ at its endpoints satisfy $xy\prec r^{\varepsilon}.$ Since $\partial D$ alternates between $R^{\varepsilon}R^{\varepsilon}$-arcs and such intervals, it follows that $\partial D$ goes through $4gn$ oriented $RR^{*}$-arcs for some $n\geq 1$. Since no piece boundary goes positively through such an oriented $RR^{*}$-arc,
@@ D. Wick product of Newton polynomials and traceless-tensors: the Gross-Taylor formula
for any pair of partitions $\lambda,\mu$ with [^51] Mind that it is true for all $\mu\in\mathbb{Yb}.$ $|\lambda|\leq N.$ Here $(Z_{i})_{i}$ are i.i.d. normal, centred, complex random variables of variance $i$ and so
@@
For $M\leq N,$ the family $(p_{[\lambda,\mu]})_{\lambda,\mu\in\mathbb{Yb},|\lambda|+|\mu|\leq M}$ is the unique orthogonal family of $L^{2}(U(N))$ with
@@
for all $\lambda,\tilde{\lambda},\mu,\tilde{\mu}\in\mathbb{Yb}$ with $|\lambda|+|\mu|,|\tilde{\lambda}|+|\tilde{\mu}|\leq M.$
Proof.
Assume first $m=0$ and recall a standard argument. For any $\sigma\in S_{n},$
@@
Denoting by $\chi(\mu)$ the value of an $S_{n}$-character $\chi$ on a conjugacy class with cycle length type $\mu,$ by orthogonality of characters $(z_{\mu}^{-\frac{1}{2}}\chi^{\lambda}(\mu))_{\lambda,\mu\vdash n}$ is an orthogonal matrix. Orthogonality of its columns together with the last display yield (295) for $m=0$ and $n\leq N.$
Assume now $m\geq 1$. For any $\alpha\times\beta\in S_{n}\times S_{m}$ by Koike-Schur-Weyl duality
@@
$\langle p_{[\alpha,\beta]},p_{[\pi,\sigma]}\rangle_{L^{2}(U(N))}=\delta_{n,n^{\prime}}\delta_{m,m^{\prime}}\sum_{\lambda\vdash n,\mu\vdash m,\ell(\lambda)+\ell(\mu)\leq N}\chi^{\lambda}(\alpha)\chi^{\mu}(\beta)\chi^{\lambda}(\pi^{-1})\chi^{\mu}(\sigma^{-1}).$
When $N\geq n+m,$ the last display equals
@@
Proof of Prop. D.1 for $|\lambda|+|\mu|\leq\frac{N}{2}$.
@@
$\prod_{i}Z_{i}^{m_{i}(\lambda)}\overline{Z}_{i}^{m_{i}(\mu)}:=Q_{\lambda,\mu}(p_{i},\overline{p}_{i},i\geq 1).$
@@
$=\delta_{\lambda,\tilde{\lambda}}\delta_{\mu,\tilde{\mu}}\prod_{i\geq 1}\mathbb{E}[|Z_{i}|^{2m_{i}(\lambda)}]\mathbb{E}[|Z|^{2m_{i}(\mu)}]m_{i}(\lambda)!m_{i}(\mu)!$
$=\delta_{\lambda,\tilde{\lambda}}\delta_{\mu,\tilde{\mu}}z_{\lambda}z_{\mu}.$
@@
(305) $p_{i}^{(a,b)}=\sum_{k=0}^{a\wedge b}(-1)^{k}{a\choose k}{b\choose k}p^{a-k}_{i}\overline{p}^{b-k}_{i}.$
@@
(306) $\mathcal{H}_{n}=\mathrm{Span}\{p_{[\lambda,\mu]}:\lambda,\mu\in\mathbb{Yb}\,\text{with}\,:|\lambda|+|\mu|=n\}.$
@@
(307) $\chi_{[\lambda,\mu]}$ $=P_{|\lambda|+|\mu|}(\chi_{\lambda}\overline{\chi}_{\mu})$
@@
where $(Z_{k})_{k\geq 1}$ are independent $\mathcal{N}_{\mathbb{C}}(0,1)$-random variables and $m_{k}(\sigma)$ stands for the number of cycles of length $k$ of a permutation $\sigma.$
@@
$\sum_{\nu\in\mathbb{Y}:\,\nu\leq\lambda\wedge\mu}(-1)^{\ell(\nu)}C_{\nu}^{[\lambda,\mu]}p_{\lambda\setminus\nu}\overline{p}_{\mu\setminus\nu}=\mathbb{E}[\prod_{i}(1-\frac{1}{N}Z_{i})^{m_{i}(\lambda)}(1+\frac{1}{N}\overline{Z}_{i})^{m_{i}(\mu)}].$

### 31 Mar 2026 v2

@@ 4.5.2. Upper-bound II: bound for shortest words representatives
when $\tau_{u},\tau_{b}$ and $\varphi=\pi_{\alpha,\alpha}$ are respectively compatible permutation diagrams and an admissible quadrangle diagram, hence with $\tau_{u},\tau_{b}\in\mathcal{D}^{0}_{\mathfrak{a},\mathfrak{b}},\alpha\in S_{\mathfrak{a},\mathfrak{b}}.$ Under this assumption, let us rename $\varphi$-arcs $\alpha$-arcs. For such a Brauer map each $\mathcal{C}_{II}$-face of $\Sigma_{\tau_{u},\pi,\tau_{b},\varphi}$ can be identified with a rectangle with two parallel $\alpha$-arcs and two parallel $\iota$-arcs labeled by a letter of $\{a_{1},b_{1},\ldots,b_{g}\}$ and its inverse. Recall that each connected component of $\partial\Sigma_{\tau_{u},\pi,\tau_{b},\varphi}$ is the concatenation of $\tau,\pi$ and $\iota$ intervals. Since $\tau_{u},\tau_{b},\pi,\mathrm{id}_{s}$ are compatible for $\mathfrak{b},$ vertices along each boundary component have same $\mathfrak{b}$-colouring say $B\in\mathcal{S}$. In the latter case call the boundary component $B$-boundary. Since $\tau_{u},\tau_{b}\in\mathcal{D}_{\mathfrak{a},\mathfrak{b}}^{0}$, reading the $\mathfrak{a}$-value of $\iota$ intervals along an oriented $B$-boundary spells out a power $p$ of a cyclic permutation of the word labeled $B$ for each cycle of length $p$ of $\tau.$ We then assign the multiplicity of the boundary component to be $p.$ Note that $p=1$ when $B\in\{W,W^{-1}\},$ since, as permutations, $\tau_{u},\tau_{b}$ act trivially on $\mathfrak{b}^{-1}(\{W,W^{-1}\}).$ Counted with their multiplicity, there are $n$ boundaries coloured $R,$ $m$ coloured $R^{-1},$ one $W$ and one $W^{-1}.$ Recalling the constraint (148), since $\alpha\in S_{\mathfrak{a},\mathfrak{b}},$ an $\alpha$-arc cannot have one endpoint in a $R$-boundary and the other in a $R^{-1}$-boundary.
@@
We will next compute and bound $\chi(\Sigma_{\tau_{u},\tau_{b},\omega,\alpha})$ cutting $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$ along some $\alpha$-arcs, extracting discs or annuli along $W$ and $W^{-1}$-boundaries of $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$, such that each component has in its boundary only $W$-intervals or only $W^{-1}$ intervals. [^28] Boundary components of pieces shall be concatenation of $\alpha$-arcs and of boundary intervals. Also, since there are no $RR^{-1}$-arcs, boundary component of pieces shall include only $R$-intervals, or only $R^{-1}$-intervals. As in [Mag25], for these ”good” components, called below pieces, the shortest length condition on $\omega$ allows to bound from above the number of $W^{\pm}R^{\pm},W^{\pm}R^{\mp}$ arcs by the ones of $RR$ or $R^{-1}R^{-1}$ arcs. [^29] Allowing components with $WW^{-1}$-arcs would break this bound. The main points here is then to prove that a variant of [Mag25] yields the correct Euler characteristic bound for the remaining components despite having possibly more types of components. The argument can be split in two, with a first part bounding the Euler characteristics, denoted $\chi_{-}$ below, of a ”neighbourhood” of the good components.
@@
Assume $X$ is a $W^{\varepsilon}R^{\varepsilon^{\prime}}$-pre-piece or a $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arc for some $\varepsilon,\varepsilon^{\prime}\in\{-1,1\}.$ Thanks to Lemma 4.16, when $X$ is a pre-piece, among pre-pieces and $WR^{*}$-arcs, $\overline{X}$ only intersects closures of $W^{\varepsilon}R^{\varepsilon^{\prime}}$-pre-pieces and $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs along the two $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs in its boundary. Consider a connected set $P^{\prime},$ union of pre-pieces and $WR^{*}$-arcs with $P^{\prime}\supset X.$ By induction on the number of pre-pieces $\overline{P}^{\prime}$ intersects, $P^{\prime}$ is either an open annulus or a disc $D$ with two distinct $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs in its boundary, and $P^{\prime}$ only intersects arcs and pre-pieces of type $W^{\varepsilon}R^{\varepsilon^{\prime}}.$
@@
The side length $e(P)$ is the number of $W^{\varepsilon}R^{\varepsilon^{\prime}}$-arcs endpoints the boundary $\partial\Sigma_{W^{\varepsilon}}\cap\overline{P}$ goes through. Going throughout $\partial\Sigma_{W^{\varepsilon}}\cap\overline{P}$ according to its orientation and reading letters for each $\alpha$-arc crossed, spells out a word or a cyclic equivalence of a word $\omega_{P}$ when $\chi(P)$ is respectively $1$ or $0.$
@@
The proof is a variation of [MP23, Lemma 5.18]. Since it is the main place where the shortest length condition plays a role, let us detail a proof for completeness.
When $\omega_{1},\omega_{2}$ are two words with letters in $\mathcal{A},$ let us write $\omega_{1}\prec_{c}\omega_{2}$ when $\omega_{1}$ is a subword of a cyclic permutation of $\omega_{2},$ and $\omega_{1}\sim_{c}\omega_{2}$ when $\omega_{1}$ is cyclic permutation of $\omega_{2}.$ Consider the word
@@
Since there is no long bloc and no long chain, $d_{m}\leq 1$ for all $m$ and whenever $d_{i}=1=d_{j}$ for some $1\leq i<j\leq l$, there is $i<k<j$ with $d_{k}=-1.$ Consequently $\sum_{i=1}^{m}d_{i}\leq 1$ and (172) follows from the last display. ∎
@@
Since all connected components of $\Sigma^{*}$ and $\Sigma_{\tau_{u},\tau_{b},\omega,\alpha}$ have non-empty boundary, they deform retract to the dual graph associated to any cell decomposition. Denote $\hat{\mathcal{G}}$ the graph dual to the $1$-skeleton of $X_{\tau_{u},\tau_{b},\omega,\alpha}$, so that vertices are labeled by $2$-cells of $X_{\tau_{u},\tau_{b},\omega,\alpha}$ and edges by $\alpha$-arcs. Let $\hat{\mathcal{G}}_{R}$ be the graph obtained from $\hat{\mathcal{G}}$ by removing all edges given by $RR$ and $R^{-1}R^{-1}$-arcs. Then $\Sigma^{*}$ deforms retract to $\hat{\mathcal{G}}_{R}$ so that
@@
The only contribution to $\chi(\Sigma^{*})$ therefore comes from joint discs. Let us split the sum (178) as
@@
Since type of boundaries alternate between $R^{*}$ and $W^{*}$ boundaries when going through $\partial D$ and crossing a $WR^{*}$-arc, $d_{WR^{*}}(D)$ is even. If $d_{WR^{*}}(D)\geq 4,$ then
@@
Since any $WR^{*}$-arc bounding a joint disc is also bounding a piece, while any piece boundary goes through exactly two distinct oriented $WR^{*}$-arcs,353535with the convention that for $WR^{*}$-arc that do not belong to any pre-piece, the boundary of the piece is exactly the arc with some orientation followed by its reverse.
@@
Denote by $\mathfrak{he}^{\prime}_{RR^{*}}$ the number of oriented $RR^{*}$-arcs through which no piece boundary goes positively. Since each oriented $RR^{*}$-arc is positively bounding at most one piece,
@@
If $D$ is a $R^{\varepsilon}$-disc, $d^{*}(D)=0$ and for each oriented interval of $\partial\Sigma_{R^{\varepsilon}}$ meeting its boundary, the $\mathfrak{a}$-values $x,y$ at its endpoints satisfy $xy\prec r^{\varepsilon}.$ Since $\partial D$ alternates between $R^{\varepsilon}R^{\varepsilon}$-arcs and such intervals, it follows that $\partial D$ goes through $4gn$ oriented $RR^{*}$-arcs for some $n\geq 1$. Since no piece boundary goes positively through such oriented $RR^{*}$-arcs,
@@ D. Wick product of Newton polynomials and traceless-tensors: the Gross-Taylor formula
for any pair of partitions $\lambda,\mu$ with [^51] Mind that it is true for all $\mu\in\mathbb{Yb}.$ $|\lambda|\leq N.$ Here $(Z_{i})_{i}$ are independent normal, centred, complex random variables of variance $i$ and so
@@
The family $(p_{[\lambda,\mu]})_{\lambda,\mu\in\mathbb{Yb},|\lambda|+|\mu|\leq N}$ is the unique orthogonal family of $L^{2}(U(N))$ with
@@
for all $\lambda,\tilde{\lambda},\mu,\tilde{\mu}\in\mathbb{Yb}$ with $|\lambda|+|\mu|,|\tilde{\lambda}|+|\tilde{\mu}|\leq N.$
Let us give an argument that builds on the one of [DS94] that we recall now for completeness.
Proof of (286).
For any $\sigma\in S_{n},$
@@
In particular (286) holds when $|\lambda|\not=|\mu|.$ Also for $n\leq N,$ denoting by $\chi(\mu)$ the value of an $S_{n}$-character $\chi$ on a conjugacy class with cycle length type $\mu,$ by orthogonality of characters $(z_{\mu}^{-\frac{1}{2}}\chi^{\lambda}(\mu))_{\lambda,\mu\vdash n}$ is an orthogonal matrix. Orthogonality of its columns together with the last display yield (295) when $|\lambda|\leq N.$
Proof of Lemma D.1.
Assume $n,m\geq 1$. For any $\alpha\times\beta\in S_{n}\times S_{m}$ by Koike-Schur-Weyl duality
@@
(298) $\langle p_{[\alpha,\beta]},p_{[\pi,\sigma]}\rangle_{L^{2}(U(N))}=\delta_{n,n^{\prime}}\delta_{m,m^{\prime}}\sum_{\lambda\vdash n,\mu\vdash m,\ell(\lambda)+\ell(\mu)\leq N}\chi^{\lambda}(\alpha)\chi^{\mu}(\beta)\chi^{\lambda}(\pi^{-1})\chi^{\mu}(\sigma^{-1}).$
When $N\geq n+m,$ since $(z_{\mu}^{-\frac{1}{2}}\chi^{\lambda}(\mu))_{\lambda,\mu\vdash n}$ is orthogonal, the last display equals
@@
Proof of Prop. D.1.
@@
$\prod_{i}Z_{i}^{m_{i}(\lambda)}\overline{Z}_{i}^{m_{i}(\mu)}:=Q_{\lambda,\mu}(Z_{i},\overline{Z}_{i},i\geq 1).$
@@
$=\delta_{\lambda,\tilde{\lambda}}\delta_{\mu,\tilde{\mu}}\prod_{i\geq 1}(i^{m_{i}(\lambda)+m_{i}(\mu)}m_{i}(\lambda)!m_{i}(\mu)!)=\delta_{\lambda,\tilde{\lambda}}\delta_{\mu,\tilde{\mu}}z_{\lambda}z_{\mu}.$
@@
(306) $p_{i}^{(a,b)}=\sum_{k=0}^{a\wedge b}(-1)^{k}k!i^{k}{a\choose k}{b\choose k}p^{a-k}_{i}\overline{p}^{b-k}_{i}.$
@@
(307) $\mathcal{H}_{n}=\mathrm{Span}\{p_{[\lambda,\mu]}:\lambda,\mu\in\mathbb{Yb}\,\text{with}\,:|\lambda|+|\mu|=n\}$
@@
(308) $\chi_{[\lambda,\mu]_{N}}$ $=P_{|\lambda|+|\mu|}(\chi_{\lambda}\overline{\chi}_{\mu})$
@@
where $(Z_{k})_{k\geq 1}$ are independent, with $Z_{k}\sim\mathcal{N}_{\mathbb{C}}(0,k)$ and $m_{k}(\sigma)$ stands for the number of cycles of length $k$ of a permutation $\sigma$ for all $k\geq 1.$
@@
$\sum_{\nu\in\mathbb{Y}:\,\nu\leq\lambda\wedge\mu}(-1)^{\ell(\nu)}C_{\nu}^{[\lambda,\mu]}p_{\lambda\setminus\nu}(\mathrm{Id}_{N})\overline{p}_{\mu\setminus\nu}(\mathrm{Id}_{N})=\mathbb{E}[\prod_{k}(N+iZ_{k})^{m_{i}(\lambda)}(N+i\overline{Z}_{k})^{m_{k}(\mu)}].$